OpenStudy (anonymous):

The function y = n + 3 describes the sequence 3, 6, 9, 12. True or False

jigglypuff314 (jigglypuff314):

false because when you plug in for n for example if n=1 then y=(1)+3=4 if n=2 y=(2)+3=5 therefore the function y=n+3 does not describe the sequence 3,6,9,12

OpenStudy (anonymous):

Thanks (:

jigglypuff314 (jigglypuff314):

your welcome :)